OpenStudy (anonymous):

s(x)=2x^2 t(x)=x^3 write the expression (t+s)(x) and (t*s)(x) and evaluate (t-s)(2).

OpenStudy (anonymous):

medal and fan

OpenStudy (kirbykirby):

\((t+s)(x)=t(x)+s(x)\) \((t*s)(x)=t(x)*s(x)\) \(t-s)(2)=t(2)-s(2)\) so for example: \((t+s)(x)+t(x)+s(x)=(2x^2)+(x^3)=2x^2+x^3\) no other simplification

OpenStudy (kirbykirby):

last line should real: \((t+s)(x)\color{red}{=}t(x)+s(x)\)

OpenStudy (anonymous):

what i don't get what your doing

OpenStudy (anonymous):

@satellite73 @Hero @zepdrix

OpenStudy (kirbykirby):

\(t(x)+s(x)\)... since \(t(x)=x^3\), and \(s(x) = 2x^2\) \(t(x) + s(x) = x^3+ 2x^2\) Maybe it's more clear now. I inverted the addition before but it doesn't matter, addition here is commutative :)

OpenStudy (anonymous):

so how do I add it into the formula

OpenStudy (kirbykirby):

that's all there is to do

OpenStudy (kirbykirby):

\((t*s)(x)=t(x)*s(x)=x^3*2x^2\), this you can simplify more using exponent laws: \(x^a*x^b=x^{a+b}\)